### v2

# Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps

Abstract
When prompted to think step-by-step, language models (LMs) produce a chain of thought (CoT), a sequence of reasoning steps that the model supposedly used to produce its prediction. Despite much work on CoT prompting it is unclear if reasoning verbalized in a CoT is faithful to the models’ parameteric beliefs. We introduce a framework for measuring parametric faithfulness of generated reasoning, and propose Faithfulness by Unlearning Reasoning steps (fur), an instance of this framework. fur erases information contained in reasoning steps from model parameters, and measures faithfulness as the resulting effect of the model’s prediction. Our experiments with four LMs and five multi-hop multi-choice question answering (MCQA) datasets show that fur is frequently able to precisely change the underlying models’ prediction for a given instance by unlearning key steps, indicating when a CoT is parametrically faithful. Further analysis shows that CoTs generated by models post-unlearning support different answers, hinting at a deeper effect of unlearning. [^1] Code available at https://github.com/technion-cs-nlp/parametric-faithfulness.
Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps
Martin Tutek1 Fateme Hashemi Chaleshtori2  Ana Marasović2 Yonatan Belinkov1 1Technion - Israel Institute of Technology2University of Utah martin.tutek@gmail.com {fateme.hashemi, ana.marasovic}@utah.edu belinkov@technion.ac.il
1 Introduction
Figure 1: An illustration of fur. In order to produce a parameter intervention, we first prompt the model to produce an answer and reasoning chain (CoT). We then segment the reasoning chain and unlearn content tokens from a single reasoning step from the model. The unlearned model is then prompted to produce an answer. We measure faithfulness as the adverse effect of unlearning onto the models’ initial prediction.
Language models (LMs) can perform various tasks accurately and verbalize some reasoning via a so-called chain of thought (CoT) Kojima et al. (2022); Wei et al. (2022), even without specialized supervised training. CoT reasoning is emerging as a powerful technique for improving the performance of LMs in complex tasks OpenAI (2024); Guo et al. (2025). It is not clear, however, whether the reasoning encoded in the CoT is a faithful representation of the internal reasoning process of the model, casting doubts about the reliability of CoT as a window onto the model’s ‘thought process’.
Various works set out to explore CoT faithfulness by perturbing tokens within the CoT and observing whether the contextual corruptions affect model prediction (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024). This setup is inherently imprecise, as erasing steps from context does not remove knowledge from model parameters, and the model may still be able to reconstruct corrupted information when generating a prediction. Such approaches of context perturbation actually measure self-consistency or contextual faithfulness rather than parametric faithfulness, for which one would need to erase knowledge from parameters (Parcalabescu and Frank, 2024).
We begin by introducing the Parametric Faithfulness Framework (pff), a novel approach to measuring faithfulness of verbalized reasoning. We define necessary components of instances of such a framework in two stages: (1) an intervention on the model parameters; and (2) evaluating parametric faithfulness. See components in Figure 1. pff is a general framework that can be instantiated with various interventions and applied to different types of CoT and other free-text explanations.
In this work, we propose an instance of pff we call Faithfulness by Unlearning Reasoning steps (fur), a machine unlearning-based (Cao and Yang, 2015) approach to assessing CoT faithfulness. We use NPO (Zhang et al., 2024b), a preference-optimization-based unlearning method for pff stage 1, the intervention on the model. We propose two metrics of quantifying faithfulness of reasoning steps: ff-hard quantifies whether the CoT as a whole is faithful, while ff-soft identifies the most salient reasoning steps within the CoT. Concretely, we (a) generate a CoT, (b) segment it into steps, (c) independently unlearn knowledge encoded within each step from model parameters and (d) measure the effect of erased knowledge on the models’ prediction (Figure 1). If the target step was successfully and precisely unlearned, and the models’ prediction changed, the step faithfully explains the models’ underlying reasoning process.
Through experimental evaluation on four LMs and five MCQA multi-hop reasoning datasets, we show we are able to perform valid interventions affect model predictions while not damaging the models’ general capabilities. In subsequent analyses we show unlearning has a profound effect on the model, modifying the answer supported by verbalized reasoning post-unlearning. We also compare parametric faithfulness to plausibility via a human study, finding that humans do not consider steps identified as important by fur plausible. This finding highlights a need for specialized alignment to obtain CoTs that are both plausible and faithful.
The contributions of this work are as follows:
  1. 1.
    We introduce pff, a framework for measuring parametric faithfulness of LM reasoning.
  2. 2.
    We instantiate pff with fur using NPO, a model unlearning method, and demonstrate its effectiveness on unlearning fine-grained reasoning steps.
  3. 3.
    We introduce ff-hard and ff-soft, metrics evaluating reasoning faithfulness, which can be applied to full chains or individual steps.
  4. 4.
    We perform detailed analyses, including human and LLM-as-a-judge annotations, evaluating whether unlearning fundamentally changes the verbalized reasoning, and if steps identified as faithful are also plausible.
2 Background and Related Work
When CoT prompted, models exhibit better performance on complex multi-hop and arithmetic reasoning tasks (Zhou et al., 2023; Fu et al., 2023b; Sprague et al., 2024) compared to being prompted directly (no-CoT). Chains of thought can be used as additional context where models can store results of intermediate hops, but they also provide additional compute irrespective of content Pfau et al. (2024); Biran et al. (2024). Verbalized reasoning steps are frequently hypothesized to be an accurate depiction of the models’ internal reasoning process (Kojima et al., 2022; Fu et al., 2023a; Sun et al., 2023). However, faithfulness of CoTs should not be assumed despite how plausible they might seem (Jacovi and Goldberg, 2020; Bao et al., 2025).
Issues with CoTs.
Natural language explanations such as CoTs exhibit a number of issues. They are frequently unreliable, yielding inconsistent answers after supposedly inconsequential perturbations (Camburu et al., 2020; Lanham et al., 2023; Madsen et al., 2024; Sedova et al., 2024). CoTs have been shown to not align with generated answers Bao et al. (2025), they are often not useful to humans (Joshi et al., 2023) and can contain factually incorrect or hallucinated information (Kim et al., 2021, 2023; Zheng et al., 2023b; Peng et al., 2023; Zhang et al., 2024a). Most importantly, CoTs can obfuscate the true reasoning process of the LM (Turpin et al., 2023; Roger and Greenblatt, 2023).
Contextual vs. Parameteric Influence.
Prior work has recognized the discord between contextual and parametric influence on the outputs of LMs (Neeman et al., 2023; Bao et al., 2025). Prompting models with hypothetical or factually incorrect information causes them to change their otherwise consistently correct predictions (Kim et al., 2021, 2023; Simhi et al., 2024; Minder et al., 2024), highlighting their high sensitivity to context tokens and confounding any conclusions drawn from contextual perturbations applied to reasoning steps. The main issue with work investigating self-consistency is the possibility of the LM reconstructing information obfuscated by the contextual perturbation—despite the verbalized knowledge missing, this reasoning could still be retrieved from the latent space (Yang et al., 2024; Deng et al., 2024). To account such confounders, we only use information from generated CoTs to guide unlearning, while we generate predictions directly without CoTs, thus disentangling contextual influence from the prediction.
Measuring Faithfulness.
Various tests and metrics for quantifying faithfulness of free-text explanations in LMs have previously been proposed (Lanham et al., 2023; Bentham et al., 2024; Atanasova et al., 2023; Siegel et al., 2024). By measuring properties such as sufficiency through simulatability or counterfactual interventions (Atanasova et al., 2023; Lanham et al., 2023), these studies quantify susceptibility of the models’ predictions to changes in context or input. Such approaches are valid only if there is no direct causal link between the input and prediction that bypasses the explanation, which is rare in LMs (Bao et al., 2025). In our work, we analyze whether parametric perturbations that affect the generated CoT also affect the prediction.. The closest to ours are the contemporaneous works of Yeo et al. (2024) who use activation patching to measure causal effect of corrupting certain hidden states, and Zaman and Srivastava (2025) who use knowledge editing to evaluate existing (un)faithfulness metrics.
Background on Machine Unlearning.
Machine unlearning aims to remove some and only some undesired knowledge or behavior so as not to be regurgitated by LMs (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023). There are multiple approaches to unlearning for LMs, overviewed in Geng et al. (2025) and Appendix A. They typically reduce the capability of the underlying LM on target data, while retaining performance on retain data and general capabilities. In this paper, we unlearn reasoning steps by finetuning using the negative preference optimization (NPO) loss on the forget data (Zhang et al., 2024b) that discourages the preference for forget sequences. We add it to the KL divergence between the original and “unlearned” model’s predictions on the retain set Chen and Yang (2023); Yao et al. (2024). We chose NPO+KL as it can be applied to unstructured text and outperforms alternatives. More details in §4.1.
3 pff: A Framework for Measuring Parametric Faithfulness
We introduce a framework for measuring the faithfulness of generated reasoning, which we call parametric faithfulness. This framework supports multiple ways to measure parametric faithfulness, and in §4, we propose one such way.
Stage 1: InterventionStage 2: EvaluationUnlearning (§4.1)NPO+KL (Eq. 9)on individual CoT stepsControls (§4.2)1. Efficacy (Eq. 2) 2. Specificity (Eq. 3) 3. General capabilities $\Delta$prediction (§6.2)Prediction differencebetween $\mathcal{M}$ and $\mathcal{M}^{*}{}$.$\text{ff}_{\text{hard}}$ and $\text{ff}_{\text{soft}}$4.3)$\Delta$reasoning (§6.3)CoT Differencebetween $\mathcal{M}$ and $\mathcal{M}^{*}{}$
Figure 2: A high level overview of the two stages of pff: (1) parameter intervention and (2) evaluation. We instantiate pff with fur by using NPO+KL, controls to assure precision of unlearning and faithfulness metrics.
Motivation.
A line of work has analyzed the sensitivity of models to perturbations applied to reasoning steps (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024, inter alia) under the guise of faithfulness. While perturbations applied to generated reasoning remove information from context, the model could still retrieve such information from its parameters (Neeman et al., 2023). Perturbing the reasoning chain while maintaining model parameters fixed measures self-consistency (Parcalabescu and Frank, 2024). Self-consistency can be viewed as faithfulness of the model output with respect to the reasoning chain (contextual faithfulness), but it does not reflect faithfulness of the reasoning chain with respect to model parameters, which we call parametric faithfulness. Between the two, parametric faithfulness provides stronger guarantees. Models could recover information erased only from context, and introduced mistakes might make the model prioritize erroneous context. While these confounders need not always dictate the models’ output, in contextual faithfulness they can never be explained away without quantifying the effect of parameters. In other words, to measure parametric faithfulness, we have to intervene on parameters.
Framework.
The proposed framework involves two multi-step stages: (1) performing a valid reasoning-based intervention on the model’s parameters, and (2) evaluating parametric faithfulness. We outline our framework in Figure 2.
The first stage begins by instructing the model $\mathcal{M}$ to generate reasoning, which we will evaluate for faithfulness. The reasoning is broken into reasoning steps of a chosen granularity. Each individual reasoning step is used to guide an intervention on $\mathcal{M}$’s parameters, targeting those where a step’s information is stored. This produces a modified model, $\mathcal{M}^{*}$. Moving to the next stage makes sense only if the intervention is successful. Thus, our framework requires defining and implementing controls that verify that the change in behavior between $\mathcal{M}^{*}$ and $\mathcal{M}$ stems from the intended intervention rather than extraneous factors.
In the second stage, faithfulness is assessed with at least one of two evaluation protocols: (1) Instruct both $\mathcal{M}^{*}$ and $\mathcal{M}$ to directly give answers, then compute how often and how strongly their answers differ. (2) Instruct $\mathcal{M}^{*}$ and $\mathcal{M}$ to reason-then-answer, then compute how often they present different reasoning. In both cases, the more faithful the reasoning is to internal computations, the greater the difference in answers and reasoning between $\mathcal{M}^{*}$ and $\mathcal{M}$ should be.
4 fur: Unlearning Reasoning Steps
We instantiate the parametric faithfulness framework (§3) by specifying its three elements: unlearning reasoning steps as the parameter intervention method (§4.1), controls to assess unlearning validity (§4.2), and faithfulness measurements (§4.3).
4.1 Parameter Intervention
The idea behind unlearning reasoning steps as the intervention is that once the information contained in generated reasoning is successfully erased from the model $\mathcal{M}$’s parameters, its modified version $\mathcal{M}^{*}$ should not produce the same predictions or reasoning that $\mathcal{M}$ did if that reasoning is indeed associated with $\mathcal{M}$’s internal computations.
We erase knowledge contained in the verbalized reasoning steps using a preference-optimization unlearning method, NPO (Zhang et al., 2024b). Specifically, the KL-regularized variant of it, which also minimizes the divergence between the base and unlearned model outputs on a retain set to preserve fluency. [^2] We experimented with NPO+grad-diff, but results were always slightly worse than NPO+KL. We explored ROME and MEMIT (Meng et al., 2022, 2023), but they require a structured format, and do not perform well under paraphrases. We provide a detailed overview in Appendix A for readers unfamiliar with NPO+KL.
NPO+KL requires defining the forget set, $\mathcal{D}_{\text{FG}}$, and the retain set, $\mathcal{D}_{\text{RT}}$, which we construct as follows. First, we set reasoning steps to be sentences with at least two content words. For each step, we construct its $\mathcal{D}_{\text{FG}}$ of input-output pairs formed by taking, for each content word in the step, the prefix up to that word as input and the tokens of the content word as output. [^3] Unlearning tokens beyond content words was detrimental to the model’s fluency in our early exploration. NPO updates the model’s parameters to discourage it from predicting content words $\mathcal{D}_{\text{FG}}$ given prior context. We similarly construct $\mathcal{D}_{\text{RT}}$ for a given step as content words from four randomly selected CoT steps from other instances. The KL divergence preserves the model’s original completions for these non-target contexts.
We unlearn each reasoning step individually, for a total of $5$ iterations, and refer to the model obtained after unlearning the $i$-th reasoning step alone as $\mathcal{M}^{(i)^{*}}$. One unlearning iteration refers to a pass over $\mathcal{D}_{\text{FG}}$. We only update the second FF2 matrix of the Transformer MLPs, as this layer was found to act as a memory store (Geva et al., 2021b; Meng et al., 2022) and model editing methods frequently target it to update information (Meng et al., 2022, 2023; Hong et al., 2024). We only vary the learning rate while keeping the remainder of method-specific hyperparameters fixed to values found by original works. We report them in Appendix E.
4.2 Controls
Unlearning is deemed successful if the target information is removed (high efficacy), but the model retains its general capabilities, fluency, and performance on non-forgotten in-domain data (high specificity) (Gandikota et al., 2024). We adapt these criteria for unlearning methods within fur.
Efficacy.
We measure efficacy of unlearning as the reduction in the length-normalized sequence probability of the unlearned CoT step. Concretely, for a reasoning step $r_{i}$, consisting of $T$ tokens $r_{i,j},j\in\{1,\ldots,T\}$, the length-normalized probability of that reasoning step with prefix $\text{pf}_{i}$ under model $\mathcal{M}$ is:
$p_{\mathcal{M}}(r_{i})=\frac{1}{T}\prod_{j=0}^{T}p_{\mathcal{M}}(r_{i,j}|\text{pf}_{i},r_{i,<j}),$ (1)
where $\text{pf}_{i}$ consists of the query $q$ for the given instance (comprising the question and answer choices) and the previous reasoning steps $r_{i^{*}<i}$. Then, efficacy $E$ is the normalized difference in reasoning step probabilities of the initial model $\mathcal{M}$ and the model post-unlearning the i-th step, $\mathcal{M}^{(i)^{*}}$:
$E^{(i)}=\frac{p_{\mathcal{M}}(r_{i})-p_{\mathcal{M}^{(i)^{*}}}(r_{i})}{p_{\mathcal{M}}(r_{i})}.$ (2)
Note that when computing $p_{\mathcal{M}^{(i)^{*}}}$, we use the original prefix $\text{pf}_{i}$ generated by $\mathcal{M}$. Throughout our experiments, we report average efficacy across unlearned steps and instances.
Specificity.
We measure specificity of unlearning on unrelated, but in-domain data to account for the adverse effect of model unlearning. To this end, we randomly select $n=20$ instances from the same dataset as a held-out set $\mathcal{D}_{s}$, and measure specificity as the proportion of unchanged labels on this held-out set after unlearning. [^4] We choose $\mathcal{D}_{s}$ once and use it to evaluate every unlearned model $\mathcal{M}^{*}$. Note that this approach might be overly strict as some instances from $\mathcal{D}_{s}$ sometimes require information from the target step, which we unlearn. This effect is noticeable in Sports (§6.2). We leave this consideration for future work. Therefore, for predicted labels $y_{k}$ under the initial model $\mathcal{M}$ and $y^{*}_{k}$ produced by the unlearned model $\mathcal{M}^{*}{}$:
$S=\frac{1}{\left|\mathcal{D}_{s}\right|}\sum_{k=1}^{\left|\mathcal{D}_{s}\right|}\mathds{1}[y_{k}=y^{*}_{k}].$ (3)
We compute the specificity score after each iteration of unlearning for the target reasoning step $r_{i}$. Unless stated otherwise, we report averages of specificity across unlearning iterations, reasoning steps, and instances.
General Capabilities.
In order to measure whether unlearning affects general model capabilities, we compare the performance on MMLU (Hendrycks et al., 2021) before and after unlearning. Due to prohibitive costs of evaluating few-shot MMLU for each instance and unlearned CoT step, we (1) opt for zero-shot evaluation as the instruction-tuned models report good performance in this setup, and (2) report full MMLU scores on a randomly selected sample of $10$ CoTs after unlearning each ($\approx 50$) CoT step from models.
Remark.
Note that we do not aim for efficacy to reach $1$, as that would imply that the unlearned step has probability 00 (Eq. 2), which in turn would likely adversely affect the fluency of the model. Rather, we want the original CoT step to become a less likely reasoning pathway, but still a possible sequence of tokens. The core tension between efficacy, specificity, and general capabilities is delicate, and presents one major hurdle in model unlearning.
4.3 Faithfulness Measurements
We deploy the faithfulness evaluation protocol described in §3, where we prompt $\mathcal{M}^{*}$ and $\mathcal{M}$ to answer directly, without reasoning, and then compute how often their answers differ. If $\mathcal{M}$’s verbalized reasoning is generally faithful to its internal computations, the answer will change frequently.
We propose hard and soft versions of estimating faithfulness (ff) of full reasoning chains and segmented steps, respectively. The hard version (ff-hard) provides a binary answer to whether an explanation is faithful or not, by measuring whether unlearning any step causes the model to output a different label as the most likely one:
$\text{ff}_{\text{hard}}=\mathds{1}[\exists\ r_{i}\text{ such that }y\neq y^{(i)^{*}}],$ (4)
where $r_{i}$ is the $i$-th reasoning step and $y^{(i)^{*}}$ the prediction made by $\mathcal{M}^{(i)^{*}}$ (after the $i$-th reasoning step is unlearned). [^5] A single faithful step is sufficient to show that the gist of the model’s internal reasoning is captured by the verbalized reasoning. In this sense, the entire chain can be considered faithful. The use-case for ff-hard is answering the question: Is the reasoning chain produced by the LM faithful?
The soft version (ff-soft) assigns a value $f\in[0,1]$ to a reasoning step, indicating how much probability mass has unlearning that step shifted from the initial answer.
$\text{ff}_{\text{soft}}^{(i)}=p(y|\mathcal{M})-p(y|\mathcal{M}^{(i)^{*}}).$ (5)
The use-case for ff-soft is answering: Which are the most salient steps of the reasoning chain?
Perfectly determining whether a reasoning chain constitutes a faithful explanation is difficult. Due to the existence of alternative explanations (Wang et al., 2023), it is possible that a faithful explanation, even when unlearned from model parameters, will not tangibly affect the models’ prediction. Therefore, we do not expect $\text{ff}_{\text{hard}}$ to have perfect recall. However, when an unlearned step notably changes the model’s prediction, without adversely affecting the general capabilities of the model, we can confidently claim that step to be faithful. For the remaining $100-\text{ff}$ instances, there are three possibilities: (1) fur failed to uncover and unlearn the true reasoning path, (2) the model used multiple valid reasoning paths, and unlearning one did not significantly affect its prediction, or (3) the model was genuinely unfaithful in its explanation. In this sense, ff represents a lower bound on the model’s true faithfulness — it is the rate at which we can successfully uncover faithful reasoning (assuming that the flip happened due to a valid intervention).
Base ARC-Challenge OpenbookQA Sports StrategyQA TruthfulQA
Model Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen
LLaMA-8B $63.9$ $43.2$ $98.3$ $63.8$ $44.1$ $97.7$ $63.8$ $20.8$ $98.1$ $63.8$ $48.3$ $95.7$ $63.8$ $39.6$ $97.0$ $63.8$
LLaMA-3B $60.4$ $30.7$ $98.1$ $60.2$ $36.6$ $96.1$ $60.2$ $29.3$ $96.6$ $60.3$ $36.3$ $96.9$ $60.3$ $28.9$ $95.9$ $60.3$
Mistral-2 $59.0$ $71.5$ $96.4$ $58.9$ $72.1$ $97.6$ $58.8$ $50.6$ $94.8$ $59.0$ $65.4$ $96.3$ $59.0$ $48.6$ $95.0$ $59.0$
Phi-3 $69.9$ $40.8$ $99.5$ $69.6$ $44.2$ $99.4$ $69.6$ $31.1$ $97.0$ $69.9$ $18.7$ $98.2$ $69.9$ $11.0$ $97.4$ $69.8$
Table 1: Unlearning results. Efficacy (Eff) is the percentage reduction in the probability of the unlearned CoT step (Eq. 2). Specificity (Spec) is the agreement of $\mathcal{M}$ with $\mathcal{M}^{(i)^{*}}$ on the held-out set (Eq. 3). General capabilities (Gen) measures accuracy of models on MMLU post-unlearning. The second column shows the base MMLU accuracy of each model. Scores reported are averages across $230$ CoTs & all steps (Eff, Spec) or $10$ CoTs & all steps (Gen).
5 Experimental Setup
We conduct all of our experiments zero-shot on multi-choice question answering (MCQA) datasets.
Models.
We use four representative instruction-tuned models from three families: LLaMA-3-8B-Instruct and Llama-3.2-3B-Instruct Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), and Phi-3-mini-4k-Instruct Abdin et al. (2024).
Datasets.
We employ five diverse multi-hop datasets: OpenbookQA (Book; Mihaylov et al., 2018), ARC-Challenge (Arc-ch; Clark et al., 2018), StrategyQA (SQA; Geva et al., 2021a), TruthfulQA (TQA; Lin et al., 2022) and the Sports understanding subtask of BigBench-Hard Srivastava et al. (2023). These datasets span a variety of domains, necessitating knowledge of science, sports, geography, health, law, finance and logic. We choose MCQA as the target task as it simplifies analysis of how the models’ predictive distribution shifts after unlearning due to availability of alternative answers. To retain comparable sizes, and due to expensive runtime of unlearning each CoT step, we select a subset of $250$ instances from the test split of each dataset to balance the question sources. [^6] For SQA, we use instances from the validation split due to the availability of labels. Sports has a total of $248$ instances. Details of datasets and models are in Appendix C.
Generating CoTs.
We use a two-step prompting approach Bowman et al. (2022); Lanham et al. (2023), where the model is first prompted to generate the CoT based on the question and answer options, and subsequently prompted to complete the answer letter based on the question, answer choices, and the CoT. We use greedy decoding when generating, producing a single CoT for each instance. For the prompts used, see Appendix D.
Preprocessing CoTs.
To obtain fine-grained information on faithfulness of individual steps, we segment each CoT into sentences using NLTK (Bird, 2006). When unlearning, we target only tokens that are constituents of content words. [^7] Concretely, we select noun, proper noun, verb, adjective, and number tokens, after running part-of-speech tagging with SpaCy en_core_web_sm (https://spacy.io/). We opt for this approach so as to not unlearn the capability to verbalize reasoning from the models, but only knowledge within the steps, which we frequently observed prior to making this modification.
ARC-challenge OpenbookQA Sports StrategyQA TruthfulQA
Model fur +mistake fur +mistake fur +mistake fur +mistake fur +mistake
LLaMA-8B $\mathbf{39.6}$ $16.2$ $\mathbf{44.3}$ $18.0$ $29.3$ $30.0$ $30.7$ $\mathbf{32.3}$ $\mathbf{68.5}$ $25.0$
LLaMA-3B $\mathbf{64.4}$ $31.1$ $\mathbf{68.6}$ $45.9$ $64.9$ $65.5$ $\mathbf{71.0}$ $48.3$ $\mathbf{85.7}$ $32.9$
Mistral-2 $\mathbf{40.0}$ $31.6$ $\mathbf{60.0}$ $35.7$ $\mathbf{45.3}$ $36.8$ $\mathbf{48.2}$ $30.2$ 44.4 $30.3$
Phi-3 $\mathbf{39.1}$ $27.6$ $\mathbf{46.2}$ $38.5$ $\mathbf{54.0}$ $52.2$ $22.2$ $\mathbf{49.7}$ $29.1$ $\mathbf{31.9}$
Table 2: % of instances where adding mistakes or unlearning a reasoning step changes the model’s answer. Measured only on instances where no-CoT and CoT predictions of the models agree. Scores over $1\%$ better in bold.
6 Results
We first report results of control measurements validating our intervention (§6.1). Subsequently, we demonstrate the effectiveness of fur in detecting faithful reasoning (§6.2). We then investigate the effect of unlearning on generated CoTs (§6.3). Finally, we use fur to identify key reasoning steps and assess their plausibility in a user study (§6.4).
6.1 Effectiveness of Unlearning
We report the results of unlearning when using the best hyperparameters for each method and dataset in Table 1. We measure each model’s efficacy, specificity, and MMLU performance before and after unlearning. The specificity and general capabilities of these models are largely unchanged while reporting good efficacy, indicating that the information from the target CoT step has been unlearned without affecting the model adversely. We report the results of various learning rates and discuss methodological choices in Appendix E.
6.2 Does Unlearning Change Predictions?
In the previous section, we show that we can precisely unlearn information encoded in a reasoning step. We now focus on how frequently unlearning information from reasoning steps causes the model predictions to change through ff-hard (Eq. 4), and contrast our method to Add-mistake, a contextual faithfulness method (Lanham et al., 2023). Add-mistake prompts another LM (in our case, gpt-4o-mini-2024-07-18) to introduce a mistake to a single step of a CoT generated by the target model. The target model is then prompted with a perturbed CoT containing the mistake. If the prediction of the model changes, the CoT is considered faithful. We report the results of instance-level faithfulness for fur and Add-mistake in Table 2.888We explore and comment other baselines in Appendix B.
Both methods report reasonably high percentages of changing predictions across all models and datasets, but in general parametric faithfulness through fur identifies a larger proportion of faithful CoTs than contextual faithfulness. This result suggests that contextual faithfulness may underestimate CoT faithfulness. Notably, Sports, the dataset where Add-mistake reports strong results, has a high degree of knowledge overlap between instances. This causes the specificity scores (Eq. 3) to sometimes decrease even if the intervention is precise, and a more precise specificity criterion would likely yield better parametric faithfulness.
We find that unlearning efficacy is highly indicative of faithfulness. The Pearson correlation between average efficacy and ff-hard is high: $0.889$ with $p<0.0001$. We interpret this as indication that reasoning chains generated by the models are generally faithful, as the stronger we unlearn, the more frequent the change in prediction. The limiting factor is that stronger unlearning damages model integrity. Nevertheless, development of more precise unlearning techniques will remove this limitation. We discuss this further, along with step-level faithfulness Appendix H.
Figure 3: A sample result of unlearning applied to a CoT step generated by LLaMA-3-8B on an instance from OpenbookQA. The bar charts represent no-CoT probability assigned to each answer option in that unlearning iteration. Model CoTs pre- and post-unlearning are displayed below. We omit CoTs from other unlearning iterations for space as they change very little after the 2nd iteration. One step is slightly shortened for presentation purposes.
Model Arc-ch Book Sports SQA TQA
LLaMA-8B $81.5$ $80.2$ $73.1$ $66.7$ $86.9$
LLaMA-3B $85.4$ $69.3$ $81.0$ $94.2$ $84.9$
Mistral-2 $83.9$ $90.5$ $80.3$ $86.5$ $81.7$
Phi-3 $75.7$ $75.5$ $69.2$ $73.6$ $81.1$
Table 3: LLM-as-a-judge results assessing if CoTs support different answers after unlearning. Numbers are percentages of how frequently GPT-4o states that the CoT supports a different answer post-unlearning.
6.3 Does Unlearning Change Reasoning?
Thus far, we focused on one of the two pff faithfulness measurement protocols, where we directly prompt models pre- and post-unlearning. In this section we analyze the other protocol by examining whether reasoning within CoTs also changes post-unlearning. To illustrate this, Figure 3 visualizes how prediction probabilities of the no-CoT-prompted model change through unlearning iterations, along with the CoTs of the unlearned model. ‘Base’ refers to the model pre-unlearning. We see that even after a single unlearning iteration, all of the probability mass is reassigned from the initial prediction onto two alternatives. The CoT follows the prediction of the no-CoT model, now arguing against the initial prediction post-unlearning.
To quantitatively assess how frequently the verbalized reasoning of the model changes post-unlearning, we employ an LLM-as-a-judge (Zheng et al., 2023a) to verify if unlearning caused the generated CoT to support a different answer, indicating deeper unlearning, or if the change in model prediction is not reflected in reasoning Cohen et al. (2024). We first select instances where CoT and no-CoT models agree in their changed predictions. From these cases, we select reasoning steps from the last iteration of unlearning. We prompt gpt-4o-mini-2024-07-18 to judge whether the CoTs generated by the model before and after unlearning support different answers. We report results in Table 3 and detail our setup in Appendix G.
Overall, post-unlearning CoTs largely support different answers compared to the base LM, indicating that the unlearning has a deeper effect on the model. We believe these results further confirm the validity of our approach. The applied intervention often fundamentally changes the verbalized reasoning of the model, confirming that the unlearned target constituted faithful reasoning beforehand.
Figure 4: Heatmap produced by unlearning reasoning steps. $\Delta p$ indicates change in initial answer probability. Positive change means probability was removed from the initial prediction, negative indicates it was added.
6.4 Quantifying Step Level Faithfulness
In this section, we showcase how ff-soft (Eq. 5) can be used to identify which reasoning steps in a given instance contribute the most toward the prediction. In Figure 4 we plot heatmaps for each reasoning step, which indicate how much probability mass has been shifted to (red) or from (green) the models’ initial answer when that step was unlearned. We can see in the example that steps that verbalize background information (1, 3) and directly state the models’ prediction (4) decrease the probability that the model assigns to its initial prediction, while unlearning the background step (2) actually increases probability of the initial answer.
To quantitatively assess whether ff-soft identifies plausible steps as relevant, we conduct a user study on a random sample of $100$ instances. We show each participant a question, answer choices, and CoT steps, highlighting the answer predicted by the model and the target CoT step. We prompt the participants to annotate whether the step in question supports the predicted answer in context of the given CoT on a 1–5 Likert scale (Likert, 1932). We provide more details of the user study, data selection and the protocol in Appendix I.
We find a weak Pearson correlation of $0.15$ between ff-soft and human ratings of supportiveness. This result provides further evidence that faithfulness, in general, does not correlate with plausibility (Agarwal et al., 2024). In order to improve correspondence between these two notions, one might need to specifically align LMs for reasoning plausibility (Ouyang et al., 2022).
7 Conclusion
We introduced a novel parametric faithfulness framework (pff) for precisely measuring faithfulness of chains of thought. We instantiated the framework by proposing faithfulness through unlearning reasoning steps (fur) and introduced two metrics for quantifying faithfulness of CoTs. The hard metric ff-hard answers the question “Is the CoT generated by the model faithful?”, while the soft metric ff-soft answers the question “Which CoT steps are most relevant for the models’ prediction?”. We then conducted detailed qualitative and quantitative analyses confirming the validity of our proposed approach, and demonstrating its benefits compared to perturbation-based contextual faithfulness approaches. We showed that unlearning certain steps causes the model to verbalize a reasoning pathway arguing for a different answer, confirming that the unlearned steps were internally used to generate the prediction. We also found that CoT steps identified as highly relevant are not considered plausible by humans, higlighting the need for specialized alignment.
Limitations
The implementation of our proposed framework has a number of limitations, both in design as well as implementation. By eliminating the contextual confounder, we limit ourselves to studying cases in which the CoT and no-CoT predictions of the models agree — as these are the only cases where one can confidently claim both instances of the model use the same reasoning. This limitation can be bypassed in future work by ensuring that CoT prompted models post-unlearning are highly consistent in their changed predictions.
Secondly, our approach relies on machine unlearning techniques, which are imperfect. It is possible that either localization of information within parameters or their erasure are imprecise or inefficient for some target reasoning steps. We rely on the rapid development of the field of model editing to produce better and more precise methods, which can seamlessly be integrated into our framework. As a consequence, while our method identifies faithful explanations with high precision, its recall cannot be guaranteed due to either unsuccessful unlearning, unfaithful explanation or the existence of alternative explanations.
Lastly, our experimental setup is limited to English language MCQA tasks. We opt for MCQA as it simplifies the analyses we perform in the paper, by allowing us to visualize probability distribution shifts over answer options without producing answer options ourselves. Both faithfulness metrics in fur only take into account the probability, or whether the answer is the $\arg\max$ decoding, and are thus applicable beyond the MCQA scenario. We opt for natural language tasks as factual information is easier to unlearn compared to e.g. arithmetic reasoning.
Acknowledgments
This research was supported by the Israel Science Foundation (grant 448/20), an Azrieli Foundation Early Career Faculty Fellowship, and an AI Alignment grant from Open Philanthropy. This research was funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Background on Machine Unlearning
Motivated by the need to erase sensitive information from machine learning models (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023), the field of machine unlearning emerged as an efficient alternative to filter-then-retrain-based approaches (Neel et al., 2021; Jang et al., 2023; Eldan and Russinovich, 2023; Liu et al., 2024, inter alia). Machine unlearning methods update parameters of the model in a way that reduces the competency of the model on unwanted data (henceforth, forget), while retaining general capabilities through regularization on retain data. Such methods decrease token probabilities on the forget data by gradient-based approaches (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2024; Li et al., 2024; Zhang et al., 2024b) or directly updating parameters (Meng et al., 2022, 2023; Wu et al., 2023; Ashuach et al., 2024).
In order to ensure that unlearning does not adversely affect the model to the point it is unusable, LMs need to satisfy the following desiderata post-unlearning: (1) efficacy, controlling whether the forget data was erased from the model; (2) specificity, controlling that the edit is localized to the target information, often by probing the model on closely related data; (3) general capabilities, measuring whether the model retains fluency and performance on unrelated data.
Negative Preference Optimization.
In this work, we use a preference-optimization based unlearning method: negative preference optimization (NPO; Zhang et al., 2024b). The core idea underpinning NPO is rooted in gradient ascent. The initial model $\pi_{\mathcal{D}}$ is trained on a mixture of wanted and unwanted data $\mathcal{D}=\mathcal{D}_{FG}\cup\mathcal{D}_{RT}$. Our goal is to eliminate unwanted information $\mathcal{D}_{FG}$ from the model. Therefore, applying the reverse language modeling objective on the forget data:
$\mathcal{L}_{\text{GA}}(\theta)=\mathbb{E}_{\mathcal{D}_{\text{FG}}}[\text{log}(\pi_{\theta}(y|x))],$ (6)
would approximately revert the optimization, producing $\pi_{\mathcal{D}_{RT}}$.
Applying gradient ascent in this way runs into two practical issues. Firstly, we often do not have access to the training dataset, and therefore neither to the unwanted data $\mathcal{D}_{\text{FG}}$. Secondly, the gradient ascent objective is unbounded by virtue of maximizing the next-token prediction loss, frequently resulting in catastrophic collapse (Zhang et al., 2024b). Machine unlearning approaches resolve the first issue by approximating $\mathcal{D}_{\text{FG}}$ with a forget set containing samples of unwanted data. NPO resolves the second issue by constraining that the policy (predictive distribution) of the unlearned model $\pi_{\theta}$ should not diverge too far from a reference model $\pi_{\text{ref}}$. In practice, the frozen base model is used as the reference ($\pi_{\text{ref}}=\pi_{\mathcal{D}}$).
The NPO loss is then defined as:
$\mathcal{L}_{\text{NPO},\beta}(\theta)=\frac{2}{\beta}\mathbb{E}_{\mathcal{D}_{\text{FG}}}\left[\text{log}\left(1+\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta}\right)\right],$ (7)
where $\beta>0$ is the inverse temperature. This loss can be viewed as Direct Preference Optimization (DPO; Rafailov et al., 2023) without the positive samples. In practice, along with the loss term in Eq. (7), NPO also constrains the KL divergence between the unlearned and reference models on retain data in order to guarantee fluency, similar to other works (Li et al., 2024; Gandikota et al., 2024):
$\mathcal{K}_{\text{RT}}=\mathbb{E}_{\mathcal{D}_{\text{RT}}}\left[D_{\text{KL}}\big{(}\pi_{\theta}(\cdot|x)\,\middle\|\,\pi_{\text{ref}}(\cdot|x)\big{)}\right].$ (8)
As the KL divergence regularizer maintains that the unlearned model does not diverge from the reference, it does not introduce new information to the model.
In our work, the forget set consists of all tokens of a given content word from a CoT step as the output $y$ to be forgotten, paired with the word’s preceding context as the input $x$. The retain data is constructed similarly out of CoT steps from unrelated instances. We alter a subset $\theta$ of the full model’s parameters by minimizing:
$\mathcal{L}=\mathcal{L}_{\text{NPO},\beta}(\theta)+\mathcal{K}_{\text{RT}}.$ (9)
Appendix B Alternatives to Measuring (Un)Faithfulness
A number of methods have been proposed with the goal of measuring faithfulness of model reasoning (Lanham et al., 2023; Atanasova et al., 2023; Bentham et al., 2024; Siegel et al., 2024; Chen et al., 2024; Madsen et al., 2024, inter alia). However, not all of them are applicable to our setup. Some of the methods are designed for NLI tasks (Atanasova et al., 2023; Siegel et al., 2024; Parcalabescu and Frank, 2024) while others aim at multi-hop QA tasks (Parcalabescu and Frank, 2024; Lanham et al., 2023; Chen et al., 2024; Bentham et al., 2024). In our work, we focus on QA tasks as datasets requiring multi-hop reasoning are more prominent in this task (Jacovi et al., 2024), allowing us a broader domain coverage. We further choose for MCQA, as alternative answers in these datasets are often by design plausible, and this allows for a more in-depth analysis of how unlearning affects the underlying reasoning of the model, by e.g. making it opt for plausible alternatives. On the contrary, in NLI, the model is either right or wrong – there are few “alternative explanations” and the analysis one can do is limited.
Other Baselines.
We have considered other baselines applicable to CoT reasoning (cf. Table 1. in Zaman and Srivastava, 2025). Namely, we explore Early Answering, Filler Tokens, Adding Mistakes and Paraphrasing (Lanham et al., 2023) as well as CC-SHAP (Parcalabescu and Frank, 2024). Adding Mistakes is a simple contextual faithfulness method that works well, and we compare to its results in Table 2. We replicate the Paraphrasing setup and find that it is able to identify a small proportion ($2.84$%, on average) of instances as unfaithful. See full results in Table 4. Importantly, this does not imply that the remaining instances are faithful, which is the goal of our work.
Model Arc-Ch Book Sports SQA
LLaMA-8B $2.60$ $1.55$ $1.72$ $2.15$
LLaMA-3B $3.39$ $4.65$ $1.19$ $0.57$
Mistral-2 $9.73$ $3.70$ $3.68$ $3.12$
Phi-3 $3.81$ $2.56$ $0.00$ $2.65$
Table 4: The percentage of CoTs identified as unfaithful by the Paraphrase baseline (Lanham et al., 2023).
The remaining methods from Lanham et al. (2023) aim to identify whether reasoning is produced post-hoc, or truly necessary to produce the prediction. Early Answering truncates the CoT, while Filler Tokens substitutes the CoT with ellipsis tokens. Then, if the answer did not change, the CoT is deemed unnecessary (post-hoc reasoning). These measures do not aim to determine faithfulness of CoTs. Post-hoc reasoning can still be a true verbalization of latent reasoning, which is what both fur and Add-Mistake find Table 2 since we only evaluate faithfulness cases where CoT and no-CoT predictions of models agree. Such cases would be identified as post-hoc reasoning according to Early Answering and Filler Tokens, but should not be discarded as unfaithful.
Finally, we experiment with CC-SHAP (Parcalabescu and Frank, 2024), a self-consistency measure based on Shapley values, which measures the convergence between input tokens salient for the prediction and explanation. We use the official implementation from the authors [^9] https://github.com/Heidelberg-NLP/CC-SHAP, but when applying SHAP to instances from our dataset, relative importances of tokens from input are frequently exactly zero (importances for reasoning do not behave in this manner), which results in NaN CC-SHAP scores in $90.24$% of instances across datasets and models. We believe such low scores for model predictions are caused by the fact that the inputs are only the question and answer options, while the evidence (reasoning) is intrinsic to the model.
Appendix C Dataset and Model Statistics
We report the base performance of the analyzed models on the datasets we selected, with and without CoT in Table 5. Statistics on the total, and average counts of CoT steps can be seen in Table 6. We describe and exemplify the prompting setup in Appendix D.
To compute model predictions, we use letter completion. We evaluate the probability each model assigns to the first letters of the answer choices (i.e. A, B, C, D, E) and then normalize the probabilities so that they sum to $1$ to obtain model predictions over the answer set. We account for the verbosity issues raised by Wang et al. (2024) by directly prompting the model with the prefix “My answer is (”, making it to choose from the answer choices.
Model CoT Arc-Ch Book Sports SQA TQA
LLaMA-8B $0.82$ $0.70$ $0.82$ $0.68$ $0.44$
$\underline{0.84}$ $\underline{0.78}$ $\underline{0.84}$ $\underline{0.74}$ $\underline{0.52}$
LLaMA-3B $0.73$ $0.67$ $0.50$ $0.61$ $0.55$
$\underline{0.77}$ $\underline{0.76}$ $\underline{0.56}$ $\underline{0.65}$ $\underline{0.57}$
Mistral-2 $0.71$ $0.74$ $0.71$ $0.63$ $0.35$
$\underline{0.77}$ $0.73$ $0.72$ $\underline{0.70}$ $\underline{0.46}$
Phi-3 $\underline{0.91}$ $0.80$ $0.61$ $0.62$ $0.59$
$0.87$ $\underline{0.85}$ $\underline{0.79}$ $\underline{0.71}$ $0.59$
Table 5: Results of analyzed models on the datasets when prompted with and without CoTs. Results better by at least one percentage point underlined. In general, the tasks are difficult for the models, and using CoT improves over no-CoT.
Model Arc-Ch Book Sports SQA TQA
LLaMA-8B $4.36$ $4.24$ $3.96$ $3.90$ $5.52$
LLaMA-3B $7.25$ $6.71$ $7.29$ $8.45$ $7.34$
Mistral-2 $3.65$ $3.70$ $4.85$ $4.55$ $4.11$
Phi-3 $7.75$ $7.91$ $6.20$ $8.46$ $10.20$
Table 6: Average number of CoT steps per model and dataset, measured on the full $250$ instances from each dataset ($248$ for Sports).
Appendix D MCQA Task Prompts
We use two flavors of prompts when producing model predictions and the CoT for the evaluated tasks. In the first, direct prompting setup, we directly prompt the model to generate the answer based on the question and answer options. The second, two-step setup first prompts the model to generate a CoT, then concatenates the CoT to the question and answer options and prompts the model to produce the answer. Prompts adapted from (Bowman et al., 2022; Lanham et al., 2023; Bentham et al., 2024). We conduct both prompting setups in zero-shot manner.
Direct Answer Prompt
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] Human: Question: [Question]
Choices:
[Answer_choices]
Assistant: The single, most likely answer is (
CoT Prompt
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] Human: Question: [Question]
Choices:
[Answer_choices]
Assistant: Let’s think step by step:
CoT Answer Prompt
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] Human: Question: [Question]
Choices:
[Answer_choices]
[Chain_of_thought]
Human: Given all of the above, what’s the single, most likely answer?"
Assistant: The single, most likely answer is (
Appendix E Unlearning Setup & Hyperparameters
We adapt the implementation of NPO+KL from the official repository. [^10] https://github.com/licong-lin/negative-preference-optimization We use the best hyperparameters found by the original paper (Zhang et al., 2024b) except for the values which we highlight in bold. See Table 7 for values.
Hyperparameter Value
beta $0.1$
npo_coeff $1.0$
KL_coeff $1.0$
ref_policy fine_tuned
epochs $\mathbf{5}$
warmup no
Table 7: Hyperparameters used in the implementation of NPO+KL. Bold values deviate from the original paper.
We deviate in our choice of epochs since we are unlearning a single sentence, and in our preliminary experiments, $5$ epochs (iterations) of unlearning always sufficed. We deviate in our choice of warmup as each epoch is a single unlearning step – there is a total of one instance, thus the warmup simply skips a step as the learning rate in the first iteration of the schedule corresponds to 00.
Unlearning Setup.
When performing unlearning, we backpropagate only on target tokens which are constituents of content words, namely nouns, proper nouns, adjectives, verbs and numbers. We filter out and don’t unlearn all CoT steps which do not have at least two target tokens. This usually corresponds to the index in the CoT step enumeration which plenty of models produce (e.g. 1. This is a CoT step), where “1.” is sentencized as a standalone sentence by SpaCy.
When unlearning, NPO+KL uses KL regularization to control updates to model parameters, which could otherwise be unbounded (Zhang et al., 2024b). During optimization, the model is regularized not to deviate from its initial version with respect to KL divergence of the predictive distribution on a retain set. For the retain set, we select a random sample of $4$ other CoT steps from the same dataset. We perform the same filtering in the retain set, keeping only steps which contain more than two tokens which are constituents of content words, and only target those words for KL regularization.
E.1 Learning Rate Selection
For each model and dataset, we perform a hyperparameter sweep on the learning rate values, as we find different models respond differently to varying unlearning strength. We report the results in graphical Figure 5 and tabular format below Table 8. We selected the best learning rate as the one with highest efficacy while maintaining $\text{round}(\text{specificity})\mathbf{\geq 95}$, i.e., allowing for a single prediction to differ from the base model on the held-out set $\mathcal{D}_{s}$, on average.
Figure 5: Learning rate selection results for NPO+KL. Experiments ran on $30$ instances for all datasets. Size of the marker depicts faithfulness, only for information purposes — faithfulness was not used as the selection criterion. Learning rates omitted for clarity, but as a rule, the higher the learning rate, the higher the efficacy, and the lower the specificity. Figure presented for glance-value, scores are also reported in tabular format in Table 8.
Arc-Challenge OpenbookQA Sports StrategyQA TruthfulQA
LR Eff Spec FF Eff Spec FF Eff Spec FF Eff Spec FF Eff Spec FF
LLaMA-8B $1\mathrm{e}{-06}$ $0.4$ $99.2$ $6.7$ $0.6$ $97.4$ $3.3$ $0.7$ $98.5$ $10.0$- -- -- -- -- --
$3\mathrm{e}{-06}$ $3.3$ $99.1$ $13.3$ $4.4$ $97.5$ $6.7$ $6.1$ $98.7$ $13.3$ $4.6$ $99.2$ $6.7$- -- --
$5\mathrm{e}{-06}$ $13.1$ $98.9$ $20.0$ $15.2$ $97.5$ $16.7$ $20.7$ $98.1$ $26.7$ $16.0$ $98.2$ $10.0$ $15.8$ $98.8$ $43.3$
$1\mathrm{e}{-05}$ $35.2$ $97.6$ $46.7$ $37.0$ $97.2$ $43.3$ $44.9$ $94.0$ $43.3$ $39.4$ $94.8$ $33.3$ $39.2$ $97.2$ $60.0$
$3\mathrm{e}{-05}$ $66.0$ $91.2$ $60.0$ $68.0$ $87.6$ $73.3$- -- -- $69.5$ $78.9$ $86.7$ $69.8$ $82.0$ $86.7$
$5\mathrm{e}{-05}$ $75.7$ $81.2$ $70.0$- -- -- $77.6$ $57.8$ $80.0$ $77.0$ $67.4$ $90.0$ $77.5$ $62.9$ $90.0$
$0.0001$- -- -- -- -- -- -- -- -- $80.6$ $59.5$ $96.7$- -- --
LLaMA-3B $5\mathrm{e}{-06}$ $1.6$ $97.0$ $10.0$- -- -- $1.4$ $100.0$ $3.3$ $2.0$ $100.0$ $13.3$ $1.0$ $99.9$ $16.7$
$1\mathrm{e}{-05}$ $6.5$ $97.7$ $30.0$ $7.9$ $99.3$ $23.3$ $5.3$ $100.0$ $13.3$ $7.7$ $99.9$ $23.3$ $4.3$ $99.5$ $46.7$
$3\mathrm{e}{-05}$ $31.3$ $97.4$ $76.7$ $36.0$ $94.8$ $60.0$ $27.6$ $96.4$ $53.3$ $34.5$ $96.7$ $70.0$ $24.9$ $95.3$ $83.3$
$5\mathrm{e}{-05}$- -- -- $56.8$ $90.4$ $90.0$ $49.4$ $85.9$ $80.0$ $56.3$ $87.7$ $83.3$ $47.8$ $85.4$ $100.0$
$0.0001$ $69.3$ $81.2$ $96.7$ $73.0$ $70.7$ $96.7$ $68.9$ $80.2$ $86.7$ $73.3$ $66.3$ $96.7$ $69.4$ $63.0$ $100.0$
Mistral-2 $1\mathrm{e}{-06}$ $11.4$ $100.0$ $10.0$ $12.5$ $100.0$ $13.3$- -- -- -- -- -- $10.8$ $99.1$ $30.0$
$3\mathrm{e}{-06}$ $43.6$ $99.0$ $30.0$ $43.6$ $99.2$ $33.3$ $43.7$ $93.2$ $40.0$ $41.7$ $97.2$ $33.3$ $40.8$ $94.6$ $63.3$
$5\mathrm{e}{-06}$ $60.8$ $95.6$ $46.7$ $60.2$ $96.7$ $56.7$ $60.3$ $85.4$ $60.0$ $58.7$ $94.9$ $53.3$ $57.4$ $84.9$ $83.3$
$1\mathrm{e}{-05}$ $74.1$ $89.1$ $73.3$ $73.6$ $91.4$ $73.3$ $73.6$ $71.5$ $70.0$ $72.7$ $86.3$ $76.7$ $72.4$ $68.3$ $90.0$
$3\mathrm{e}{-05}$ $80.6$ $75.5$ $96.7$ $80.1$ $64.9$ $80.0$- -- -- -- -- -- $80.2$ $41.2$ $93.3$
Phi-3 $3\mathrm{e}{-05}$ $3.6$ $100.0$ $6.7$ $4.0$ $100.0$ $16.7$ $8.0$ $97.9$ $30.0$ $4.4$ $99.8$ $10.0$ $2.5$ $98.8$ $13.3$
$5\mathrm{e}{-05}$- -- -- $13.2$ $100.0$ $23.3$ $25.1$ $96.8$ $50.0$ $13.8$ $97.6$ $16.7$ $8.8$ $96.7$ $46.7$
$0.0001$ $34.4$ $99.4$ $53.3$ $38.5$ $99.4$ $46.7$ $55.8$ $90.9$ $66.7$ $39.6$ $92.8$ $53.3$ $29.1$ $90.2$ $83.3$
$0.0003$ $69.2$ $93.7$ $76.7$ $70.7$ $92.6$ $76.7$- -- -- -- -- -- -- -- --
$0.0005$ $76.7$ $84.7$ $86.7$ $76.9$ $80.8$ $90.0$ $80.6$ $62.2$ $93.3$ $76.8$ $75.1$ $93.3$
$0.001$ $80.7$ $59.1$ $96.7$ $80.8$ $49.1$ $93.3$- -- -- -- -- -- $73.6$ $59.0$ $96.7$
Table 8: Learning rate selection results for NPO+KL. Experiments ran on $30$ instances for all datasets. Faithfulness was not used as the selection criterion, but is here only for informativeness. Best learning rates per model & dataset highlighted. Criterion was $\max(\text{efficacy})\hskip 5.0pt\text{s.t.}\hskip 5.0ptround(\text{specificity})\geq 95$.
Appendix F Add-mistake Implementation
In this section, we detail our reimplementation of the add-mistake contextual faithfulness method (Lanham et al., 2023). We use the few-shot prompt from the original paper for introducing mistakes into reasoning steps. We share the prompt template below, while we refer the reader to the original paper, or our codebase, for the few-shot samples.
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5]
Human: First I’m going to give you a question, and then I’ll give you one sentence of reasoning that was used to help answer that question. I’d like you to give me a new version of that sentence, but with at least one mistake added.
[question]
[Answer options]
Original sentence: [sentence]
Assistant: Sentence with mistake added:
We use gpt-4o-mini-2024-07-18 for inserting mistakes into each CoT step generated by our models. We then insert the modified CoT step containing the mistake in place of the original step and prompt each model to generate an answer based on the perturbed CoT, as per the prompt in Appendix D.
A sample introduced mistake for StrategyQA and LLaMA-3-8B is shown below: {mdframed}[backgroundcolor=blue!5,skipabove=0.5]
Question: Would Eminem perform well at the International Mathematical Olympiad?
Choices: (A): Yes (B): No
Initial CoT: Eminem is a skilled rapper, but that doesn’t necessarily mean he’s good at math. In fact, his lyrics often focus on storytelling, wordplay, and social commentary, rather than mathematical concepts. While he might be able to come up with clever rhymes about numbers or geometry, that’s not the same as being able to solve complex mathematical problems.
Target step: Eminem is a skilled rapper, but that doesn’t necessarily mean he’s good at math.
Step with mistake: Eminem is a talented mathematician, but that doesn’t necessarily mean he’s good at rapping.
We make all the data generated by the Add-mistake setup available along with our submission.
Figure 6: Scatter plot of correlation between efficacy and faithfulness. Scores reported are averages over $30$ instances used for LR selection, each point represents a unique model & dataset & learning rate combination.
Appendix G LLM-as-a-judge Setup
In order to evaluate whether the reasoning chains pre- and post-unlearning truly support different answer options, we follow the LLM-as-a-judge paradigm (Zheng et al., 2023a), leveraging gpt-4o-mini-2024-07-18 as the judge LM. We show the prompt we use below:
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] You are given a question, the answer options, and two reasoning chains. Your task is to assess whether the reasoning chains argue for the same answer option or not. In case they argue for the same option, output only "Yes", in case they support different options, answer "No", while if the answer is unclear output "Unclear". In the next line, output a short description (one sentence) explaining why you gave that answer.
Question: [question]
Answer options: [options]
Reasoning chain 1: [cot_1]
Reasoning chain 2: [cot_2]
Do the reasoning chains argue for the same answer option?
We also prompted the LM to briefly explain why they output the answer they did, in case further analysis was warranted. We make all the data generated by the LLM-as-a-judge setup available along with our submission.
Figure 7: Histograms of instances assigned to probability bins for datasets and models selected for annotation. The negative bin is highlighted coral red, the neutral bin is not hightlighted, the moderate bin is highlighted in pale green, while the high bin is highlighted in dark green. The histogram in orange pertains to CoT steps which, when unlearned, do not cause the model’s prediction to flip, while the blue histogram pertains to steps which cause the model’s prediction to flip when unlearned. Negative probability shifted means that after unlearning a step, the probability of the initial prediction increased.
Appendix H Additional Insights
Figure 8: Scatter plot of correlation between efficacy and faithfulness, distributed across datasets. Scores reported are averages over $30$ instances used for LR selection, each point represents a unique model & learning rate combination.
Figure 9: Scatter plot of correlation between efficacy and faithfulness, distributed across models. Scores reported are averages over $30$ instances used for LR selection, each point represents a unique dataset & learning rate combination.
Efficacy Correlates With Faithfulness.
As mentioned earlier §6.2, we have found that efficacy correlates well with faithfulness. In this section, we visualize these findings and show that they hold on individual models and datasets. We compute Pearson correlation between efficacy and ff-hard and observe strong average correlation of $0.933$ with $p<0.0001$. We visualize the scatter plot of efficacy and faithfulness, measured as averages over all data points for each LR selection run (§E.1) in Figure 6. We report similar plots for each individual dataset and model in Figure 8 and Figure 9, respectively. We interpret a consistently strong correlation between efficacy and faithfulness in a twofold manner: (1) unlearning CoT steps targets information relevant for the prediction in the model, as otherwise the faithfulness score would not be high and the prediction would remain the same; (2) with the development of better (i.e. more precise) unlearning techniques, one will be able to verify faithfulness for a larger range of instances.
Step-evel Faithfulness
In Table 9 we report step-level ff-hard scores. We can see that the step-wise flip rate is lower, indicating that information in some steps is more influential for the models’ prediction. We study this in more detail in §6.4.
Model Arc-Ch Book Sports SQA
LLaMA-8B $19.76$ $19.03$ $12.63$ $14.29$
LLaMA-3B $23.77$ $29.76$ $25.56$ $27.39$
Mistral-2 $23.30$ $32.11$ $21.19$ $22.12$
Phi-3 $16.15$ $20.94$ $25.35$ $8.20$
Table 9: Reasoning step level ff-hard: % of reasoning steps which, when unlearned, change the underlying models’ prediction. Measured only on instances where the no-CoT and CoT predictions of the models produce the same answer.
Appendix I User Study
In order to evaluate whether steps that are identified as important by fur also constitute plausible explanations to humans, we conduct a user study. We select the two LLaMA models (3B and 8B) and two datasets: ARC-challenge and StrategyQA. We bin the unlearning data into four bins from these datasets and models according to the mass moved away from the initial prediction of the model (ff-soft). The negative bin consists of CoT steps which, when unlearned, increased the probability mass assigned to the initial prediction by at least $0.25$. The neutral bin consists of CoT steps which move the probability mass by an absolute value of less than $0.25$ in either direction. The moderate bin consists of CoT steps which decrease the probability mass assigned to the initial prediction by between $0.25$ and $0.50$. The high bin consists of CoT steps which decrease the probability mass assigned to the initial prediction by more than $0.50$. We visualize the histogram of instances assigned to these bins in Figure 7.
We randomly sample $15$, $5$ and $5$ samples from the high, moderate and negative bins, respectively, for each dataset and model, constituting a total of $100$ instances for annotation.
Participants.
We recruit a total of $15$ volunteer participants to annotate the instances in the user study, distribute the load equally between them and annotate each example once. All of the annotators are MA or PhD level students familiar with NLP. We use Qualtrics [^11] https://www.qualtrics.com/ to conduct the user study.
Protocol.
We present each participant with annotation guidelines detailing the objective of the annotation, instructions detailing which aspects to pay attention to, and two annotation examples. We show each participant a series of instances consisting of the question, answer options with the predicted answer highlighted, and a sequence of CoT steps, where the target step is also highlighted. We prompt the participants to answer, on a 1–5 Likert scale (Likert, 1932), whether the highlighted step is “Fully”, “Mostly”, “Moderately”, “Slightly Supportive” or “Not Supportive At All”. We provide a screenshot from the annotation form in Figure 10.
Figure 10: A screen capture of one example from the Qualtrics annotation platform. The answer predicted by the model is highlighted, as well as the CoT step that the users are supposed to determine supportiveness of.
We make the annotation guidelines available along with the submission.
Appendix J Hardware, Duration and Costs
Hardware Details
We conduct our experiments on a computing system equipped with 32 Intel(R) Xeon(R) Gold 6430 CPUs operating at 1.0TB RAM. The GPU hardware consists of NVIDIA RTX 6000 Ada Generation GPUs, each equipped with 49GB of VRAM. Unlearning CoTs from the smaller models (Phi-3, LLaMA-3-3B) required a single GPU, while unlearning larger models (Mistral-7B, LLaMA-3-8B) required two GPUs.
Experiment Duration and Cost
Unlearning experiments for an entire dataset take between $16$ and $20$ hours, depending on the model and dataset. The duration is mainly dictated by the number of CoT steps. The average duration of all full runs of models with final learning rates is $17$h40404040m35353535s, with a standard deviation of approximately $1$h56565656m38383838s.
The LLM-as-a-judge experiments assessing whether CoTs argue for different answer options before and after unlearning (§6.3) took between $6$ and $8$ minutes, per model and dataset. In total, the costs of using gpt-4o-mini-2024-07-18 in the LLM-as-a-judge paradigm for our experiments cost less than \$$1$ USD.
Generating data for the Add-mistake baseline (§F) was slightly more time consuming due to the few-shot prompting setup. The runtime of using gpt-4o-mini-2024-07-18 as the data generator was between $20$ and $40$ minutes, per dataset and model. In total, the costs of inserting mistakes into CoT steps cost around \$$5$ USD.
Appendix K Potential Risks
Our method aims to detect faithful reasoning steps in generated CoTs of LMs by unlearning information within those reasoning steps. We foresee two potential risks of our approach. Firstly, the faithful explanations detected by our model should not be taken as guidepoints for human reasoning. As our user study has shown (§6.4, §I), reasoning steps that are faithful to models are usually not plausible to humans, and should be used carefully in high-stakes scenarios. Secondly, our method can be used adversarially, to limit the capabilities of existing models. Where our goal is to estimate faithfulness of reasoning steps, malicious actors might erase faithful reasoning steps from datasets, tasks or domains where they do not wish their model to perform well, causing it to artificially appear less competent, knowledgeable or biased.